DIRC Operations Weekly Report

Period: Wednesday 2007/7/26 - Wednesday 2007/8/01


Oncall: Rolf Andreassen
Fast Monitoring: Rolf Andreassen, Jose Benitez

Contents:

    PEP Activities     BaBar Activities     DIRC activities     Backgrounds     Scheduled activities     Pending Work     Links

Luminosities this week

PEP peak lumi: ~1.02 x 1034 (last week: ~1.0 x 1034)

PEP delivery: 4.64 fb-1 (last week: 2.97 fb-1)

BaBar integrated lumi:4.5 fb-1 (last week: 2.81 fb-1)

Run 6 total delivery: 69.09 fb-1

Run 6 total BaBar recorded: 66.45 fb-1 total, including 4.18 fb-1 off-peak data

PEP activities/issues this week

  • Jul 27-29 : BCM commissioning, teething troubles.
  • Aug 01: Install helium lines to look for vacuum leak.

    BaBar activities/issues this week

  • DIRC activities/issues this week

  • Jul 30: Alarm thresholds changed.
  • Jul 31: Neutron counters' power supply tripped - Jerry fixed this.

    See DIRC logbook for details

    Lumi and Background this week

    Scheduled activities

    Pending Work

  • Unplug Christmas tree in Sector 5.

    Manpower situation

    Justine, Jose and Rolf rotating backup & on-call commissioning duties.

    Links